# ReGLA: Refining Gated Linear Attention

### v2

@@ Abstract
Peng Lu1 Ivan Kobyzev2 $\dagger$ Mehdi Rezagholizadeh3* Boxing Chen2 Philippe Langlais1 $\dagger$ 1DIRO, Université de Montréal 2Huawei Noah’s Ark Lab  3Advanced Micro Devices, Inc. peng.lu@umontreal.caivan.kobyzev@huawei.commehdi.rezagholizadeh@amd.com boxing.chen@huawei.comfelipe@iro.umontreal.ca
@@ 3.1 Feature Maps
We posit that the arbitrary value of the inner product of broadly feature map functions could induce training instability in the forward propagating, which cannot be addressed by adding a normalization layer Qin et al. (2022) after the implicit inner product calculation. The issue comes from the unbounded value of the inner product of the features. We address this problem by using the normalized exponential feature mapping function. Assume that $x\in\mathbb{R}^{d\times L}$ a sequence of vectors of length $L$ and hidden size $d$. Define the corresponding query and key feature map as follows [^1] Note that there are alternative functions to replace $\max(\cdot)$. For example, $\log\sum\exp(\cdot)$ also ensures that the resulting inner-product remains bounded and this is equivalent to use a softmax function as the feature mapping.:
@@
where $i\in[1,d]$ is the index of dimension and $l\in[1,L]$ is the order index of input element, ${\bm{W}}_{q}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}_{k}\in\mathbb{R}^{d\times d}$ are learnable model parameters. Notice that the dot product of features is always bounded: $0<\phi_{q}({\bm{x}})^{\top}\phi_{k}({\bm{x}})\leq d$, where $d$ is the dimension of keys and queries. Indeed, each component of a vector of the form $\exp(z-\max(z))$ is bounded between 00 and $1$, so the dot product can be upper-bounded by $d$ and lower-bounded by 00.
@@ 3.2 Normalization
ReGLA $\exp({\bm{x}}-\max({\bm{x}}))$/ ✓ ${\bm{S}}_{t}={\bm{F}}_{t}\odot{\bm{S}}_{t-1}+{\bm{v}}_{t}\phi({\bm{k}}_{t})^{\top},{\bm{F}}_{t}=\left((\mathbb{1}-{\bm{r}}_{t})\odot{\bm{g}}_{t}^{2}+{\bm{r}}_{t}\odot(\mathbb{1}-(\mathbb{1}-{\bm{g}}_{t})^{2})\right)\mathbb{1}^{\top},$

### v3

# Peng Lu1 Ivan Kobyzev2 Mehdi Rezagholizadeh3* Boxing Chen2 Philippe Langlais1 $\dagger$

@@ Abstract
Peng Lu1 Ivan Kobyzev2 Mehdi Rezagholizadeh3* Boxing Chen2 Philippe Langlais1 $\dagger$ 1DIRO, Université de Montréal 2Huawei Noah’s Ark Lab  3Advanced Micro Devices, Inc. peng.lu@umontreal.caivan.kobyzev@huawei.commehdi.rezagholizadeh@amd.com boxing.chen@huawei.comfelipe@iro.umontreal.ca
@@ 3.1 Feature Maps
We posit that the arbitrary value of the inner product of broadly feature map functions could induce training instability in the forward propagating, which cannot be addressed by adding a normalization layer Qin et al. (2022) after the implicit inner product calculation. The issue comes from the unbounded value of the inner product of the features. We address this problem by using the normalized exponential feature mapping function. Assume that $x\in\mathbb{R}^{d\times L}$ - a sequence of vectors of length $L$ and hidden size $d$. Define the corresponding query and key feature map as follows [^1] Note that there are alternative functions to replace $\max(\cdot)$. For example, $\log\sum\exp(\cdot)$ also ensures that the resulting inner-product remains bounded and this is equivalent to use a softmax function as the feature mapping.:
@@
where $i\in[1,d]$ is the index of dimension and $l\in[1,L]$ is the order index of input element, ${\bm{W}}_{q}\in\mathbb{R}^{d\times d}$ and ${\bm{W}}_{k}\in\mathbb{R}^{d\times d}$ are learnable model parameters. Notice that the dot product of features is always bounded: $0<\phi_{q}({\bm{x}})^{\top}\phi_{k}({\bm{x}})\leq d$, where $d$ is the dimension of keys and queries. Indeed, each component of a vector of the form $\exp(z-\max(z))$ is bounded between 0 and $1$, so the dot product can be upper-bounded by $d$ and lower-bounded by 0.
@@ 3.2 Normalization
ReGLA $\exp({\bm{x}}-\max({\bm{x}}))$ ✗/ ✓ ${\bm{S}}_{t}={\bm{F}}_{t}\odot{\bm{S}}_{t-1}+{\bm{v}}_{t}\phi({\bm{k}}_{t})^{\top},{\bm{F}}_{t}=\left((\mathbb{1}-{\bm{r}}_{t})\odot{\bm{g}}_{t}^{2}+{\bm{r}}_{t}\odot(\mathbb{1}-(\mathbb{1}-{\bm{g}}_{t})^{2})\right)\mathbb{1}^{\top},$